# Local Surface Parameterizations via Geodesic Splines

### v1

@@ 1. Introduction
However, most of the approaches developed for computing local parameterizations only work on meshes, or more generally, graphs. This allows them to leverage existing tools such as discrete differential operators (Sharp et al., 2019; Herholz and Alexa, 2019) or Dijkstra’s algorithm (Schmidt et al., 2006; Schmidt, 2013; Melvær and Reimers, 2012), but ties output quality to sampling or mesh quality. The sources of geometry in modern applications are quite varied (from large artist-generated meshes, to LiDAR-scanned point clouds, to hand-crafted implicits, and even neural network-encoded implicit functions (Takikawa et al., 2021; Müller et al., 2022)) and for many of these, it can be difficult to extract the high-quality samples and topological information required to apply previous methods. Meshes, for instance, may have low-quality triangles that are unsuitable for finite element-based computation; point clouds lack explicit topology altogether and may contain noise from the acquisition procedure; and implicit functions lack both explicit geometry and topology, with possibly noisy isosurfaces in the case of neural implicits. Many methods exist to address these problems, from remeshing (Botsch and Kobbelt, 2004), to surface reconstruction (Lorensen and Cline, 1987; Ju et al., 2002; Chen and Zhang, 2021; Chen et al., 2022; Sellán et al., 2023), and even sampling the surface (Witkin and Heckbert, 1994; De Goes et al., 2012), but these methods are quite disparate, and as such there is no unified way to generate local parameterizations for all these types of geometry.
@@
We propose a technique to avoid such preprocessing by only requiring a very narrow interface from the input geometry: a signed implicit function, along with a projection function derived from this implicit function. It is typically straightforward to compute this implicit representation: for example, obtaining an implicit function from meshes is simply a matter of computing the unsigned minimum distance and signing it with the winding number (Jacobson et al., 2013; Barill et al., 2018). Our algorithm proceeds in two stages: first tracing tangent plane paths from a prescribed origin, and then fusing them into a continuous local map using a spline surface. Tracing geodesics along surfaces via projection is a simple yet surprisingly delicate procedure (Fig. 3), so we also present a curvature-sensitive substepping procedure to take robust steps in high-curvature regions, as well as an inter-curve smoothing method that greatly reduces map artifacts created by large variations in curvature across the surface. To our knowledge, our method is the first to develop robust radial tracing for local parameterizations. By exclusively relying on pointwise implicit function queries rather than global surface operators, we obtain an output-sensitive method (i.e., where computational work scales in proportion to the size of the map) that produces local parameterizations meeting or exceeding the quality of previous methods, particularly on surfaces other than high-quality triangle meshes. We demonstrate the efficacy of our approach on a wide variety of geometric representations and regions on these geometries.
@@ 2. Related Work
Computing geodesics. Several methods have been developed to compute geodesics on meshes, both exact and approximate (Polthier and Schmies, 2006; Surazhsky et al., 2005; Kimmel, 1998; Crane et al., 2017). This problem comes in many flavours, from tracing out geodesic paths, to finding globally shortest geodesics between two given points, and even to finding all geodesics from one point to the rest of a set of points — see Crane et al. (2020) for a more comprehensive survey. Many of these methods can operate in a purely intrinsic setting due to them having access to an underlying mesh, which allows them to avoid an extrinsic projection-like procedure, but as such they are limited in the kinds of geometry they can operate on. On the other hand, our work removes the underlying mesh assumption and works on a broader set of geometry, though we note that we focus only on the geodesic tracing problem to build our parameterizations, rather than the full set of geodesic problems that have been studied. Some recent work also explores regularized geodesic distances (Edelstein et al., 2023), but this is primarily restricted to meshes as well. While there has been some prior work on computing geodesics on implicit surfaces which also leverages a projection operator derived from the implicit function (Pedersen, 1995), it requires an underlying blue noise global sampling of the surface to warm start the procedure, which our method avoids. Projection can also behave poorly in high-curvature regions (Fig. 3), which is why many previous methods prefer intrinsic mesh-based operations; however, we utilize a substepping procedure that preemptively avoids difficult projection scenarios, making it a feasible building block for our parameterization.
@@ 3. Method
[…]
@@ 3.1. Radial Tracing
We wish to find a local parameterization about a point $\mathbf{p}\in\mathcal{M}$, which we denote by $q_{\mathbf{p}}\colon\mathcal{T}_{\mathbf{p}}\mathcal{M}\to\mathcal{M}$ (we do not write it as $\exp_{\mathbf{p}}$ since, as we will describe later, we do not always want an exact exponential map). To start, we set $q_{\mathbf{p}}(\mathbf{0})=\mathbf{p}$, and then we numerically integrate $m$ equally spaced geodesic-like radial curves radiating from $\mathbf{p}$, with initial tangent directions defined extrinsically in $\mathbb{R}^{3}$. Using $T_{\mathbf{p}}\mathcal{M}$ to represent the tangent plane at $\mathbf{p}$ to $\mathcal{M}$ (i.e., the embedding of $\mathcal{T}_{\mathbf{p}}\mathcal{M}$ in $\mathbb{R}^{3}$), these tangent directions are denoted by $\mathbf{t}_{i,0}\in T_{\mathbf{p}}\mathcal{M}$, $10\leq i\leq m-$ 1. We perform a single integration step in direction $i$ by moving along $\mathbf{t}_{i,0}$ for a prescribed step size $h$, and projecting to $\mathcal{M}$:
@@ 3.2. Substepping
Although $h$ is small, projecting a straight line from a general $T_{\mathbf{q}_{i,j}}\mathcal{M}$ onto $\mathcal{M}$ can still behave poorly far away from the plane origin: its projection image can shrink significantly to almost zero length in regions of very high curvature, and can even split near the medial axis (Fig. 3). Although smoothing the normal field prevents image splitting, a deformed projection image can significantly degrade the quality of the final parameterization. Therefore, we sometimes need to take substeps smaller than $h$ to ensure that each full step has a geodesic length of roughly $h$, which is a technique commonly used when numerically integrating differential equations in order to obtain higher accuracy (Dahlquist and Björck, 2008).
@@ 3.3. Holonomy Smoothing
Ideally, we want no variation in acceleration (i.e., constant acceleration) so the radial curves spread out uniformly. The surfaces satisfying this property are those with constant Gaussian curvature, so we will try to direct the curves so that they behave as if they were on a constant curvature surface. On such surfaces, we expect the total Gaussian curvature in the wedges between curves to be equal, where wedges are defined as $\Delta_{i,j}=[\mathbf{p},\ldots,\mathbf{q}_{i,j},\mathbf{q}_{i+1,j},\ldots,\mathbf{p}]$ (see inset) where each point is connected by a geodesic to the next point in the sequence and $i$ indices are modulo $m$. Rather than directly integrating the curvature within this loop, which is difficult to do on an implicitly defined surface, we will instead use holonomy, which is defined as the amount of rotation induced by parallel transport along a closed loop on the surface, and equivalently the total Gaussian curvature contained in that loop on the surface. This way, we can convert an integration problem into a parallel transport problem, which we are well-equipped to solve. Since we want to find in-plane rotations, we represent the rotations by angles $\theta_{i,j}$ in each $\mathcal{T}_{\mathbf{q}_{i,j}}\mathcal{M}$ and solve for these angles. Denoting the holonomy in a wedge by $R(\Delta_{i,j})$, we smooth out the wedge holonomies by minimizing a Dirichlet-like energy to obtain rotations $\bm{\Theta}_{j}=[\theta_{0,j},\theta_{1,j},\ldots,\theta_{m-1,j}]$:
@@
We measure total wedge holonomy by decomposing the wedge loop into strips $S_{i,k}=[\mathbf{q}_{i,k-1},\mathbf{q}_{i,k},\mathbf{q}_{i+1,k},\mathbf{q}_{i+1,k-1},\mathbf{q}_{i,k-1}]$ (see inset), where each point is connected by a geodesic just like in the definition of $\Delta_{i,j}$, and then adding the holonomies of these strips. Starting at $\mathbf{q}_{i,k-1}$, our goal is to transport $\mathbf{t}_{i,k-1}$ around $S_{i,k}$. and to simplify the result, we will set each $\mathbf{t}_{p,q}$ to have an angle of 0 in its corresponding tangent space. By construction, transporting along the geodesic from $\mathbf{q}_{i,k-1}$ to $\mathbf{q}_{i,k}$ does not change the angle, and the same is true from $\mathbf{q}_{i+1,k}$ to $\mathbf{q}_{i+1,k-1}$, so the main challenge is transporting from $\mathbf{q}_{i,k}$ to $\mathbf{q}_{i+1,k}$ ($\mathbf{q}_{i+1,k-1}$ to $\mathbf{q}_{i,k-1}$ is similar). When the distance between $\mathbf{q}_{i,k}$ and $\mathbf{q}_{i+1,k}$ is small, we can approximately transport vectors from $\mathcal{T}_{\mathbf{q}_{i,k}}\mathcal{M}$ to $\mathcal{T}_{\mathbf{q}_{i+1,k}}\mathcal{M}$ using the smallest rotation that aligns $T_{\mathbf{q}_{i,k}}\mathcal{M}$ with $T_{\mathbf{q}_{i+1,k}}\mathcal{M}$, in the same manner by which we transported tangent vectors along short geodesic segments during tracing. Then, we can find the change in angle (induced by a change of basis between tangent spaces) by measuring the angle between the transported $\mathbf{t}_{i,k}$ (denoted $\hat{\mathbf{t}}_{i,k}$) and $\mathbf{t}_{i+1,k}$ in $\mathcal{T}_{\mathbf{q}_{i,k}}\mathcal{M}$, which is just the signed angle from $\mathbf{t}_{i+1,k}$ to $\hat{\mathbf{t}}_{i,k}$ in $\mathbb{R}^{3}$. When $\theta_{i,k}=\theta_{i+1,k}=0$ (i.e., before smoothing), we denote this angle by $\phi_{i,k}$, so in the general non-zero case, we have an angle of $\theta_{i,k}+\phi_{i,k}-\theta_{i+1,k}$ (Fig. 5). The change in angle from $\mathbf{q}_{i+1,k-1}$ to $\mathbf{q}_{i,k-1}$ has a similar expression but negated since we travel in the opposite direction with respect to the indices; thus we have
@@
Surprisingly, this expression does not directly depend on the angles obtained from previous steps.
@@ 3.4. Map Interpolation
After tracing out the points, the next step is to interpolate these points to form a continuous map from the planar disc $D_{R}\subset\mathcal{T}_{\mathbf{p}}\mathcal{M}$ ($R=nh$) to the surface: $q_{\mathbf{p}}\colon D_{R}\to\mathcal{M}$. Each input point $\mathbf{x}=(x,y)\in D_{R}$ can be expressed in polar coordinates $(r,\theta)$, where $r\in[0,R]$, and from there, we evaluate the surface as a spline in polar coordinates. First, we interpolate the isolines $\Psi$ using periodic cubic splines as functions of angle. We then evaluate all these splines at $\theta$ and $\theta+\pi$ (i.e., the opposite angle) to get $12n+$ 1 unique points that sample two opposite radial curves connected at the map origin, which we can connect using a natural cubic spline $\tilde{\gamma}_{\theta}$ and evaluate at $r$.

### v2

# Local Surface Parameterizations via Smoothed Geodesic Splines

@@ Abstract.
copyright: acmlicensedjournal: TOGjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3767323ccs: Computing methodologies Parametric curve and surface modelsccs: Computing methodologies Shape analysis
@@ 1. Introduction
However, most of the approaches developed for computing local parameterizations only work on meshes, or more generally, graphs. This allows them to leverage existing tools such as discrete differential operators (Sharp et al., 2019; Herholz and Alexa, 2019) or Dijkstra’s algorithm (Schmidt et al., 2006; Schmidt, 2013; Melvær and Reimers, 2012), but ties output quality to sampling or mesh quality. The sources of geometric representations in modern applications are quite varied (from large artist-generated meshes, to LiDAR-scanned point clouds, to hand-crafted implicits, and even neural network-encoded implicit functions (Takikawa et al., 2021; Müller et al., 2022)) and for many of these, it can be difficult to extract the high-quality samples and topological information required to apply previous methods. Meshes, for instance, may have low-quality triangles that are unsuitable for finite element-based computation; point clouds lack explicit topology altogether and may contain noise from the acquisition procedure; and implicit functions lack both explicit geometry and topology, with possibly noisy isosurfaces in the case of neural implicits. Many methods exist to address these problems, from remeshing (Botsch and Kobbelt, 2004), to surface reconstruction (Lorensen and Cline, 1987; Ju et al., 2002; Chen and Zhang, 2021; Chen et al., 2022; Sellán et al., 2023), and even sampling the surface (Witkin and Heckbert, 1994; De Goes et al., 2012), but these methods are quite disparate, and as such there is no unified way to generate local parameterizations for all these types of representations.
@@
We propose a technique to avoid such preprocessing by only requiring a very narrow interface from the input geometry: a signed implicit function, along with a projection function derived from this implicit function. It is typically straightforward to compute this implicit representation: for example, obtaining an implicit function from meshes is simply a matter of computing the unsigned minimum distance and signing it with the winding number (Jacobson et al., 2013; Barill et al., 2018). Our algorithm proceeds in two stages: first tracing geodesic-like paths from a prescribed origin, and then fusing them into a continuous local map using a spline surface. Tracing geodesics along surfaces via projection is a simple yet surprisingly delicate procedure (Fig. 3), so we also present a curvature-sensitive substepping procedure that avoids path-splitting steps in high-curvature regions, as well as an inter-curve smoothing method that greatly reduces map artifacts created by large variations in curvature across the surface. To our knowledge, our method is the first to develop a radial tracing method for local parameterizations that explicitly addresses both of these issues. By exclusively relying on pointwise implicit function queries rather than global surface operators, we obtain an output-sensitive method (i.e., where computational work scales in proportion to the size of the map) that produces local parameterizations meeting or exceeding the quality of previous methods, particularly on surfaces other than high-quality triangle meshes. We demonstrate the efficacy of our approach on a wide variety of geometric representations and regions on these geometries.
@@ 2. Related Work
Computing geodesics. Several methods have been developed to compute geodesics on meshes, both exact and approximate (Polthier and Schmies, 2006; Surazhsky et al., 2005; Kimmel, 1998; Crane et al., 2017). This problem comes in many flavours, from tracing out geodesic paths, to finding globally shortest geodesics between two given points, and even to finding all geodesics from one point to the rest of a set of points — see Crane et al. (2020) for a more comprehensive survey. Many of these methods can operate in a purely intrinsic setting due to them having access to an underlying mesh, which allows them to avoid an extrinsic projection-like procedure, but as such they are limited in the kinds of representations they can operate on. On the other hand, our work removes the underlying mesh assumption and works on a broader set of representations, though we note that we focus only on the geodesic tracing problem to build our parameterizations, rather than the full set of geodesic problems that have been studied. Some recent work also explores regularized geodesic distances (Edelstein et al., 2023), but this is primarily restricted to meshes as well. While there has been some prior work on computing geodesics on implicit surfaces which also leverages a projection operator derived from the implicit function (Pedersen, 1995), it requires an underlying blue noise global sampling of the surface to warm start the procedure, which our method avoids. Projection can also behave poorly in high-curvature regions (Fig. 3), which is why many previous methods prefer intrinsic mesh-based operations; however, we utilize a substepping procedure that preemptively avoids difficult projection scenarios, making it a feasible building block for our parameterization.
@@ 3. Method
Key to the generality of our approach is the limited interface it requires from input geometry. We assume that we are only given an implicit function $f(\mathbf{x})$ which contains the surface of interest in its zero isosurface $\mathcal{M}$. From $f$, we can obtain a gradient $\nabla f(\mathbf{x})$, a normal $\mathbf{n}(\mathbf{x})=\frac{\nabla f(\mathbf{x})}{\|f(\mathbf{x})\|}$ (defined only on $\mathcal{M}$), and a projection operator $\bm{\pi}(\mathbf{x})$ onto $\mathcal{M}$. We define $\bm{\pi}$ based on the general projection procedure described in Atzmon et al. (2019), which generalizes signed distance projection through an iterative root-finding procedure $\overline{\mathbf{x}}_{i+1}=\overline{\mathbf{x}}_{i}-f(\mathbf{x})\frac{\nabla f(\mathbf{x})}{\|\nabla f(\mathbf{x})\|^{2}}$, $\overline{\mathbf{x}}_{0}=\mathbf{x}$. It is worth noting that, in order to have a well-defined $\mathbf{n}$ (and hence well-defined tangent planes), 0 must not be a local extremum of $f$; in particular, unsigned distance functions do not work unless they are offset to have a different zero isosurface. Also, unless otherwise stated, we use a smoothed gradient $\widetilde{\nabla f}(\mathbf{x})=\left(\int_{B_{\mathbf{x},\epsilon}}\nabla f(\mathbf{y})d\mathbf{y}\right)/|B_{\mathbf{x},\epsilon}|$ instead of $\nabla f(\mathbf{x})$ to define normals, tangent planes, and projections, where $B_{\mathbf{x},\epsilon}$ is the ball of radius $\epsilon$ centered at $\mathbf{x}$ and $|B_{\mathbf{x},\epsilon}|$ is its volume, to remove $C^{1}$ discontinuities in $f$. This essentially turns the medial axis into regions of rapidly changing gradients, which becomes useful for detecing large accumulated curvature (Section 3.2).
@@ 3.1. Radial Tracing
We wish to find a local parameterization about a point $\mathbf{p}\in\mathcal{M}$, which we denote by $q_{\mathbf{p}}\colon\mathcal{T}_{\mathbf{p}}\mathcal{M}\to\mathcal{M}$ (we do not write it as $\exp_{\mathbf{p}}$ since, as we will describe later, we do not always want an exact exponential map). To start, we set $q_{\mathbf{p}}(\mathbf{0})=\mathbf{p}$, and then we numerically integrate $m$ equally spaced geodesic-like radial curves radiating from $\mathbf{p}$, with initial tangent directions defined extrinsically in $\mathbb{R}^{3}$. Using $T_{\mathbf{p}}\mathcal{M}$ to represent the tangent plane at $\mathbf{p}$ to $\mathcal{M}$ (i.e., the embedding of $\mathcal{T}_{\mathbf{p}}\mathcal{M}$ in $\mathbb{R}^{3}$), these tangent directions are denoted by $\mathbf{t}_{i,0}\in T_{\mathbf{p}}\mathcal{M}$, $0\leq i\leq m-1$. We perform a single integration step in direction $i$ by moving along $\mathbf{t}_{i,0}$ for a prescribed step size $h$, and projecting to $\mathcal{M}$:
@@ 3.2. Substepping
Although $h$ is small, projecting a straight line from a general $T_{\mathbf{q}_{i,j}}\mathcal{M}$ onto $\mathcal{M}$ can still behave poorly far away from the plane origin: its projection image can shrink significantly to almost zero length in regions of very high curvature, and can even split near the medial axis (Fig. 3). Since $\mathbf{n}$ is a smoothed normal, it already prevents image splitting, but can instead lead to deformed projection images, which significantly degrade the quality of the final parameterization. Therefore, we sometimes need to take substeps smaller than $h$ to ensure that each full step has a geodesic length of roughly $h$, which is a technique commonly used when numerically integrating differential equations in order to obtain higher accuracy (Dahlquist and Björck, 2008).
@@ 3.3. Holonomy Smoothing
We model our optimization on the case of constant acceleration, to encourage the curves to spread out as uniformly as possible. The surfaces satisfying this ideal property are those with constant Gaussian curvature, so the optimization procedure should encourage the curves to behave as if they were on a constant curvature surface. On such surfaces, we expect the total Gaussian curvature in the wedges between curves to be equal, where wedges are defined as $\Delta_{i,j}=[\mathbf{p},\ldots,\mathbf{q}_{i,j},\mathbf{q}_{i+1,j},\ldots,\mathbf{p}]$ (see inset) where each point is connected by a geodesic to the next point in the sequence and $i$ indices are modulo $m$. Rather than directly integrating the curvature within this loop, which is difficult to do on an implicitly defined surface, we will instead use holonomy, which is defined as the amount of rotation induced by parallel transport along a closed loop on the surface, and equivalently the total Gaussian curvature contained in that loop on the surface. This way, we can convert an integration problem into a parallel transport problem, which we are well-equipped to solve. Since we want to find in-plane rotations, we represent the rotations by angles $\theta_{i,j}$ in each $\mathcal{T}_{\mathbf{q}_{i,j}}\mathcal{M}$ and solve for these angles. (More precisely, we are now measuring the holonomy of a slightly modified $\Delta_{i,j}$ where infinitesimally small geodesics are extended from each $\mathbf{q}_{i,j}$ after rotating each geodesic tangent vector by $\theta_{i,j}$, though with a slight abuse of notation we will refer to both the infinitesimally extended wedge and the original wedge by $\Delta_{i,j}$, as we are exclusively concerned with the infinitesimally extended version.) Denoting the holonomy in a wedge by $R(\Delta_{i,j})$, we smooth out the wedge holonomies by minimizing a Dirichlet-like energy to obtain rotations $\bm{\Theta}_{j}=[\theta_{0,j},\theta_{1,j},\ldots,\theta_{m-1,j}]$:
@@
We measure total wedge holonomy by decomposing the wedge loop into strips $S_{i,k}=[\mathbf{q}_{i,k-1},\mathbf{q}_{i,k},\mathbf{q}_{i+1,k},\mathbf{q}_{i+1,k-1},\mathbf{q}_{i,k-1}]$ (see inset), where each point is connected by a geodesic just like in the definition of $\Delta_{i,j}$, and then adding the holonomies of these strips. Starting at $\mathbf{q}_{i,k-1}$, our goal is to transport $\mathbf{t}_{i,k-1}$ around $S_{i,k}$. and to simplify the result, we will set each $\mathbf{t}_{p,q}$ to have an angle of 0 in its corresponding tangent space. By construction, transporting along the geodesic from $\mathbf{q}_{i,k-1}$ to $\mathbf{q}_{i,k}$ does not change the angle, and the same is true from $\mathbf{q}_{i+1,k}$ to $\mathbf{q}_{i+1,k-1}$, so the main challenge is transporting from $\mathbf{q}_{i,k}$ to $\mathbf{q}_{i+1,k}$ ($\mathbf{q}_{i+1,k-1}$ to $\mathbf{q}_{i,k-1}$ is similar). When the distance between $\mathbf{q}_{i,k}$ and $\mathbf{q}_{i+1,k}$ is small, we can approximately transport vectors from $\mathcal{T}_{\mathbf{q}_{i,k}}\mathcal{M}$ to $\mathcal{T}_{\mathbf{q}_{i+1,k}}\mathcal{M}$ using the smallest rotation that aligns $T_{\mathbf{q}_{i,k}}\mathcal{M}$ with $T_{\mathbf{q}_{i+1,k}}\mathcal{M}$, in the same manner by which we transported tangent vectors along short geodesic segments during tracing. Then, we can find the change in angle (induced by a change of basis between tangent spaces) by measuring the angle between the transported $\mathbf{t}_{i,k}$ (denoted $\hat{\mathbf{t}}_{i,k}$) and $\mathbf{t}_{i+1,k}$ in $\mathcal{T}_{\mathbf{q}_{i,k}}\mathcal{M}$, which is just the signed angle from $\mathbf{t}_{i+1,k}$ to $\hat{\mathbf{t}}_{i,k}$ in $\mathbb{R}^{3}$. When $\theta_{i,k}=\theta_{i+1,k}=0$ (i.e., before smoothing), we denote this angle by $\phi_{i,k}$ (and all such angles at step $k$ by $\Phi_{k}$), so in the general non-zero case, we have an angle of $\theta_{i,k}+\phi_{i,k}-\theta_{i+1,k}$ (Fig. 5). The change in angle from $\mathbf{q}_{i+1,k-1}$ to $\mathbf{q}_{i,k-1}$ has a similar expression but negated since we travel in the opposite direction with respect to the indices; thus we have
@@
Surprisingly, this expression does not directly depend on the angles obtained from previous steps, but in fact the change in angle terms $\Phi_{j}$ implicitly encode this information, because they are derived from the tangent directions that were obtained from the previous step’s smoothing angles $\Theta_{j-1}$; in turn, $\Theta_{j-1}$ depends on the previous step’s $\Phi_{j-1}$, which depend on $\Theta_{j-2}$, and so on. Unrolling the dependency chain reveals that $\Phi_{j}$ contains information from all previous holonomy smoothing steps.
@@ 3.4. Map Interpolation
After tracing out the points, the next step is to interpolate these points to form a continuous map from the planar disc $D_{R}\subset\mathcal{T}_{\mathbf{p}}\mathcal{M}$ ($R=nh$) to the surface: $q_{\mathbf{p}}\colon D_{R}\to\mathcal{M}$. Each input point $\mathbf{x}=(x,y)\in D_{R}$ can be expressed in polar coordinates $(r,\theta)$, where $r\in[0,R]$, and from there, we evaluate the surface as a spline in polar coordinates. First, we interpolate the isolines $\Psi$ using periodic cubic splines as functions of angle. We then evaluate all these splines at $\theta$ and $\theta+\pi$ (i.e., the opposite angle) to get $2n+1$ unique points that sample two opposite radial curves connected at the map origin, which we can connect using a natural cubic spline $\tilde{\gamma}_{\theta}$ and evaluate at $r$.